OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. If p is the number of pounds of jelly beans Terri purchases, what expression could be used to represent the number of pounds of chocolate drops she purchases?

OpenStudy (solomonzelman):

J=Jelly C=Chocolate J = 0.98 per pound C = 0.67 per pound 2.1 *x=1.56 x is the number of pound of J and C

OpenStudy (solomonzelman):

0.98+0.98=1.96 0.98+0.67=165 So it is approximately 1 lb J and 1.1 lb C

OpenStudy (solomonzelman):

I meant 1.65

OpenStudy (solomonzelman):

It's can't be J+J b/c it's more than 1.65

OpenStudy (solomonzelman):

So the best is 1J+1.1C, which is most equal to 1.65

OpenStudy (anonymous):

Thanks! :3

OpenStudy (solomonzelman):

TY, I finally saw the approach to this. YW!